3. Problems and Applications 93 of course), but they leave some time for their favorite Frances and Latasha are roommates. They spend most of their time studying ( activities: making pizza and brewing root beer. Frances takes 3 hours to brew a gallon of root beer and 2 hours to make a pizza. Latasha takes 7 hours to brew a gallon of root beer and 5 hours to make a pizza Francess opportunity cost of making a pizza is of root beer, and Latashas opportunity cost of making a pizza of root beer. has an absolute advantage in making pizza, andh has a comparative advantage in making pizza If Frances and Latasha trade foods with each other will trade away pizza in exchange for root beer. The price of pizza can be expressed in terms of gallons of root beer. The highest price at which pizza can be traded that would maks both roommates better off is is Y of root beer, and the lowest price that makes both roommates better off Y of root beer per pizza
Answer #1

(a) Frances' opportunity cost of pizza is 0.67 (= 2/3) root beer, and Latasha's opportunity cost of pizza is 0.71 (= 5/7) root beer.

(b) Frances has absolute advantage in pizza (since she takes less time to make 1 pizza than Latasha: 2 < 5) and Frances has comparative advantage in pizza (since her opportunity cost of 1 pizza is lower than that for Latasha: 0.67 < 0.71).

(c) If they trade food, Frances will trade away pizza.

(d) Trade is beneficial when relative price (RP) lies between opportunity cost of traders. So acceptable range of RP is:

0.67 root beer < RP of 1 pizza < 0.71 root beer. Therefore,

Highest price is 0.71 (= 5/7) root beer and lowest price is 0.67 (= 2/3) root beer.